OpenStudy (anonymous):

how do you divide 7700 by 12,000? Thanks, Please show your work.

OpenStudy (inkyvoyd):

7700/12000 right?

OpenStudy (anonymous):

By using a calculator..

OpenStudy (inkyvoyd):

\(\Huge \frac{7700}{12000}\)

OpenStudy (inkyvoyd):

First we cancel out the zeros \(\Huge \frac{77}{120}\)

OpenStudy (inkyvoyd):

Next, we do long division.

OpenStudy (anonymous):

@waterineyes if ur going to be sarcastic, I dont need ur help.

OpenStudy (inkyvoyd):

This is a bit hard to do with a mouse, but I'll try my hardest.

OpenStudy (anonymous):

\[= \frac{77 \cancel{0} \cancel {0}}{120 \cancel{0} \cancel {0}} = \frac{77}{120}\]

OpenStudy (inkyvoyd):

Before we go on, however, we need to make sure that this fraction is completely reduced. Is it?

OpenStudy (anonymous):

See, I am just kidding @RuSSian1909 ..

OpenStudy (anonymous):

@inkyvoyd I think its reduced

OpenStudy (anonymous):

@waterineyes otay

OpenStudy (inkyvoyd):

Yes, that's correct @RuSSian1909 . Are you looking for a decimal approximation, or just a reduced fraction?

OpenStudy (anonymous):

decimal

OpenStudy (inkyvoyd):

Okay. To how many digits?

OpenStudy (anonymous):

just like 4 will be fine

OpenStudy (inkyvoyd):

Okay. Just a side question, is this a homework question? Or do you just have trouble with long division?

OpenStudy (anonymous):

i have trouble with long division and its for a state test

OpenStudy (inkyvoyd):